OpenStudy (anonymous):

"solve for x and y: (y-99)/(x-19) = (x-19)/y."

OpenStudy (stamp):

\[\frac{y-99}{x-19}=\frac{x-19}{y}\]

OpenStudy (anonymous):

okthenwhat